1.) Find the interest paid on a 3-year lease for a $25750 car if the car's...

1.)

Find the interest paid on a 3-year lease for a $25750 car if the car's residual value is $17591 and the lease has an annual interest rate of 4%.

Round your answer to the nearest dollar.

2.)

Find the interest paid on a 3-year lease for a $26145 car if the car depreciates at an annual rate of 11% and the lease has an annual interest rate of 5.3%.

Round your answer to the nearest dollar.
Answer #1

1.

Formula used:

2.

Formulas used:

Note:

Residual value :

Annual depreciation = $26,145 * 11% =2875.95.

Depreciation for 3 Years = 2875.95 * 3 = 8627.85

Residual value = $26,145 - $8,627.85 = $17,517.